OpenStudy (anonymous):

In an 8 gram solution of water and alcohol, the ratio by mass of water to alcohol is 3 to 1 . If 12 grams of a solution consisting of 2 grams water for each gram of alcohol is added to the 8 grams solution, what fraction by mass of the new solution is alcohol?

hero (hero):

I'm just guessing but I think this is the system 6/8 + 2/8 = 8/8 8/12 + 4/12 = 12/12 Since 2/8 + 4/12 represents the fraction by mass of the alcohol then 2/8 + 4/12 = 6/24 + 8/24 = 14/24 = 7/12

hero (hero):

So 7/12

OpenStudy (anonymous):

ah, i got it. thank you

hero (hero):

How do you know it is correct?

OpenStudy (anonymous):

because i don't understand clearly the question, so when i see your solution, i can figure what kind of problem is. thank for asking to check my understanding :)

hero (hero):

I don't even know it is correct. I told you I was just guessing. Find out what the correct answer is and let me know.